OpenStudy (anonymous):

Determine which ordered pairs are part of the solutions set for the inequality. y is less than or equal to x + 2. A. (-1, 1) B. (0, 3) C. (2, 4) D. (3, 2) True or False: The point (0, 7) is a solution to y is less than or equal to negative 3 x + 7. True False

OpenStudy (anonymous):

Using the equation see if y is equal to or less than x+2 when you plug the numbers in.

OpenStudy (texaschic101):

nelson is correct...you are going to have to sub your answer choices into the inequality and see which one is true. y <= x + 2 A. (-1,1) x = -1 and y = 1 now sub 1 <= -1 + 2 1 <= 1 (this one is correct) B. (0,3) x = 0 and y = 3 3 <= 0 + 2 3 <= 2 (incorrect) C. (2,4) x = 2 and y = 4 4 <= 2 + 2 4 <= 4 (this one is correct) D. (3,2) x = 3 and y = 2 2 <= 3 + 2 2 <= 5 (this is correct) your solutions can be A, B, and D. ======================= y <= -3x + 7 (0,7) x = 0 and y = 7 now sub 7 <= -3(0) + 7 7 <= 0 + 7 7 <= 7 TRUE. ....7 is equal to 7

OpenStudy (anonymous):

omg thanks so much

OpenStudy (texaschic101):

sure thing :)